# Energy-Efficient Channel Decoding for Wireless Federated Learning: Convergence Analysis and Adaptive Design

### v2

@@ II-C LDPC Decoding
In this work, we use LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL. In LDPC, the iterative decoding process continues until $Q$ decoding iterations, denoted as the maximum number of decoding iterations, are completed [30]. Such iterative LDPC decoding processing leads to high energy consumption. To save the decoding energy, early termination techniques are usually adopted. Specifically, if the decoded codewords can fully pass the parity check, the decoding process will stop even if not all the $Q$ decoding iterations has been completed[31]. The early termination scheme can save decoding iterations without compromising the BER performance. The energy consumption of LDPC decoding depends on the target BER and thus the pre-defined maximum number of decoding iterations. As a result, to minimize the energy consumption, it is essential to determine the maximum tolerable BER by FL systems, considering its unique statistical learning nature.
@@ III-A Effect of BER on the Convergence Rate of Wireless FL
With error-free communication, the convergence rate of FL with non-convex loss functions is proved as $\mathcal{O}(1/\sqrt{T})$ [20], where $T$ is the total number of SGD iterations. In this subsection, we investigate how BER affects the convergence rate of FL and determine the maximum admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$.
Unlike FL with analog communication[32, 33, 34, 35, 36, 37, 38], quantifying the impact of BER on FL with digital communication is not straightforward. In particular, we need to consider different cases of bit errors and analyze the impact of different bit errors, e.g., most significant bits and least significant bits.
To conduct the theoretical analysis, we take the following assumptions that are widely adopted in the literature.
@@
(Bounded Local and Global Gradient Dissimilarity [39]) The dissimilarity between the local and global gradient is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
@@
(Independent Bit Errors [40]) Each bit in the bitstream of the decoded model parameters encounters error independently.
With the $N$-bit representation, each model parameter has $12^{N}-$ 1 possible distorted patterns. However, our empirical results in Fig. 4 show that it is sufficient to focus on cases with at most one bit error of each model parameter. In particular, Fig. 4 compares the actual model error with the analytical model error expression in (7) assuming at most 1-bit error, which match closely with various BER. Therefore, we impose the following assumption to simplify our analysis.
@@
(At Most 1-bit Error of each Model Parameter) There is at most 1-bit error in the $N$-bit representation of each model parameter.
Based on Assumptions 4 and 5, we are able to establish the relationship between the decoded model at a client and the broadcasted global model in the following lemma.
@@
Denote $\tilde{\mathbf{w}}$ as the decoded model of $\mathbf{w}$. Based on Assumptions 4 and 5, the bias of each dimension in $\mathbf{w}$ is given as
$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right).$ (6)
The mean square error caused by bit errors is given as
$\mathbb{E}[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}]=\frac{d\left(4^{N}-1\right)}{3\left(2^{N}-1\right)^{2}}\cdot b(1-b)^{N-1}M^{2},$ (7)
where $b$ is the BER.
@@
Please refer to Appendix V-A. ∎
Figure 3: The difference between received noisy model and broadcasted perfect global model.
(a) BER=$10^{-2}$
(b) BER=$10^{-3}$
(c) BER=$10^{-4}$
Figure 4: Validation of the model error formula by the experiment of vanilla-CNN on Fashion-MNIST dataset. The experimental setting is in Section IV-A. In every communication round, we compute the actual model error between the noisy DNN model and perfect model, and compare with the value from the analytical function in (7). It is shown that the value from analytical function is very close to the actual value, for different BER levels and model range in different communication rounds.
After understanding the effect of BER, we analyze the convergence property of wireless FL with communication errors.
@@
With Assumptions 1-5 and the learning rate $\eta$ satisfies
$1-\frac{n+L}{n}L\eta-2L^{2}\eta^{2}E(E-1)\geq 0,$ (8)
@@
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}\\ &+\frac{L(L+2)d\hat{M}^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}\\ &+\frac{d\hat{M}^{2}}{\eta RE}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(n+1)(E-1)}{n},\end{split}$ (9)
where $f^{*}$ is the minimum global training loss, $\hat{M}$ is the maximum range of DNN weights over all communication rounds, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
@@
Please refer to Appendix V-B. ∎
@@
It is worthy noting that the first term on the RHS of (9) captures the distance between the initial and minimum global training loss. The second and third terms indicates how bit errors in each communication round affect learning convergence. The last two terms stem from non-IID data among clients and the random mini-batch sampling.
We next simplify (9) by applying $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, which satisfies (8) with a sufficient large value of $R$.
@@
With Assumptions 1-5 and a sufficient large value of $R$ such that $\eta=\frac{1}{L\sqrt{RE}}$ satisfies (8), (9) can be simplified as follows:
$\begin{split}\frac{1}{T}&\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}\\ &+\frac{C_{2}\sum_{r=0}^{R-1}(b_{r}(1-b_{r})^{N-1})^{2}}{\sqrt{T}}+\frac{C_{3}}{\sqrt{T}}+\frac{C_{4}}{T},\end{split}$ (10)
where $C_{0}\triangleq 2L(\mathbb{E}f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{L^{2}(L+2)dM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}}$, $C_{2}\triangleq LdM^{2}$, $C_{3}\triangleq\frac{(L+1)(\sigma_{L}^{2}+\sigma_{G}^{2})}{n}$, $C_{4}\triangleq\frac{L^{2}\sigma_{L}^{2}(n+1)(E-1)}{n}$.
According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round such that the RHS of (10) reduces to zero at a rate of $\mathcal{O}(1/\sqrt{T})$. Since the explicit relationship between $b_{r}$ and $r$ is difficult to obtain, we propose a simple BER regulation method in Corollary 1. As such, wireless FL systems can achieve the same convergence rate as those with error-free communication.
@@
the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss funcation at a rate of $\mathcal{O}(1/\sqrt{T})$.
@@
When $b_{r}\propto\frac{1}{(r+1)^{2}}\triangleq\frac{\vartheta}{(r+1)^{2}}$, to prove the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
$\lim_{T\rightarrow\infty}\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}=0$ (12)
due to the fact that $0\leq\left(b_{r}(1-b_{r})^{N-1}\right)^{2}\leq b_{r}(1-b_{r})^{N-1}\leq 1$ for $0\leq b_{r}\leq 1$. Specifically, since
@@
Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/(r+1)^{2})$ can guarantee wireless FL converges at a rate of $\mathcal{O}(1/\sqrt{T})$. The proposed BER regulation rule achieves an improving communication quality with communication rounds. This is consistent with the solution developed in [36], which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication round.
III-B BER vs. Number of Iterations
The analysis in Section III-A indicates that the convergence rate of wireless FL with error-free communication can be achieved if the BER is properly controlled. In this subsection, we analyze the relationship between BER and the maximum number of LDPC decoding iterations, based on which, we develop an adaptive decoding scheme to maintain the same convergence rate.
1 Input: The initial global model $\mathbf{w}_{0}$
2 Output: The converged global model $\mathbf{w}_{R}$
@@
4       for Client index $k=1,\cdots,n$ do
5             $b_{r}\propto(1/(r+1)^{2})$ with $b_{0},b_{R-1}$
6             $Q=f^{-1}(b_{r})$
7             for $decoding.iter\leq Q$ do
8                  one iteration of LDPC decoding
9                   if Early termination is satisfied then
10                        break
@@ 6             $Q=f^{-1}(b_{r})$
14            output: $\mathbf{\tilde{w}}_{r}$
@@ 17                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta\nabla f_{k}(\mathbf{w}^{k}_{r,e};\xi^{k}_{r,e})$
19            $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$
@@
21      $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{n}\sum^{n}_{k=1}\Delta\mathbf{w}_{r}^{k}$
@@
Algorithm 1 Regulating LDPC decoder in FL
It has been shown in [41] that for message-passing-based channel decoders to achieve a fraction of 1-$\epsilon$ of the channel capacity over a binary erasure channel, the relationship between the decoding complexity (denoted as $\chi_{D}$) and BER (denoted as $b$) can be modeled as follows:
@@
In the next subsection, we develop an adaptive scheme to control the maximum number of decoding iterations for clients in a wireless FL system based on Corollary 1 and (16).
III-C Decoding Iteration Control

### v3

@@ II-C LDPC Decoding
In this paper, we take the LDPC codes as an example to illustrate how adaptive channel decoding can help reduce the energy consumption of wireless FL. In LDPC decoders, the iterative decoding process continues until $Q$ decoding iterations, which is termed as the maximum number of decoding iterations, are completed [24]. Such iterative LDPC decoding processing leads to high energy consumption. To save the decoding energy, early termination techniques are usually adopted. Specifically, if the decoded codewords can fully pass the parity check, the decoding process will stop even if not all the $Q$ decoding iterations have been completed[25]. Therefore, early termination can save some decoding iterations without compromising the BER performance. The energy consumption of LDPC decoding depends on the target BER and thus the pre-defined maximum number of decoding iterations. As a result, to minimize the decoding energy consumption, it is essential to determine suitable BERs over communication rounds that can be tolerated by wireless FL systems, considering its unique statistical learning nature.
@@ III-A Effect of BER on the Convergence Rate of Wireless FL
With error-free communication, the convergence rate of FL with non-convex loss functions was proved as $\mathcal{O}(1/\sqrt{T})$ [16], where $T=RE$ is the total number of local training steps. In this subsection, we investigate how BER affects the convergence rate of FL and determine the admissible BER for maintaining the convergence rate of $\mathcal{O}(1/\sqrt{T})$. Unlike FL with analog communication where model distortion is considered as additive noise[26, 27, 28, 29, 30, 31], quantifying the impact of BER on FL with digital communication is not straightforward. In particular, we need to consider different cases of bit errors and analyze the impact of different bit errors, e.g., most significant bits and least significant bits.
To conduct the theoretical analysis, we take the following assumptions that are widely adopted in literature.
@@
(Bounded Local and Global Gradient Dissimilarity [32]) The dissimilarity between the local and global gradients is bounded, i.e., $||\nabla f_{k}(\mathbf{w})-\nabla f(\mathbf{w})||^{2}\leq\sigma_{G}^{2}$.
@@
(Independent Bit Errors [33]) Each bit in the bitstream of the model parameters encounters error (i.e., bit flipping) independently.
With the $N$-bit representation, each model parameter has $12^{N}-$ 1 possible distorted patterns. However, our empirical results in Fig. 4 show that it is sufficient to focus on cases with at most one bit error of each model parameter. In particular, Fig. 4 compares the actual model error with the analytical model error expression to be presented in (7) assuming at most 1-bit error, which match closely under various BER conditions. Therefore, we also impose the following assumption to simplify our analysis.
@@
There is at most 1-bit error in the $N$-bit representation of each model parameter.
Based on Assumptions 4 and 5, we are able to establish the relationship between the decoded model at a client and the broadcast global model in the following lemma.
@@
Denote $\tilde{\mathbf{w}}$ as the decoded model of $\mathbf{w}$. Based on Assumptions 4 and 5, the $d$-th dimension of $\tilde{\mathbf{w}}$ is given as
$\tilde{w}_{d}=w_{d}+\delta\left(w_{d}\right),$ (6)
where $w_{d}$ is the $d$-th dimension of $\mathbf{w}$ and $\delta\left(w_{d}\right)$ is the distortion term. The mean square model error caused by bit errors is given as
$\mathbb{E}[||\mathbf{\tilde{w}}-\mathbf{w}||^{2}|\mathbf{w}]=\frac{D\left(4^{N}-1\right)}{3\left(2^{N}-1\right)^{2}}\cdot b(1-b)^{N-1}M_{\mathbf{w}}^{2},$ (7)
where $b$ is the BER and $M_{\mathbf{w}}\triangleq\max(\mathbf{w})-\min(\mathbf{w})$.
@@
Please refer to Appendix A. ∎
Figure 3: Distortion between the decoded and broadcast global model. Different error bits, e.g., the most significant bit (MSB) and least significant bit (LSB), lead to different model distortion.
(a) $\text{BER}=10^{-2}$.
(b) $\text{BER}=10^{-3}$.
(c) $\text{BER}=10^{-4}$.
Figure 4: Validation of the model error analysis in (7) by experiments on the Fashion-MNIST dataset. The experimental setting is detailed in Section IV-A. In each communication round, we compute the actual model error between the decoded and broadcast global model on a selected client, and compare with the value calculated by the RHS in (7) with $\mathbf{w}$ setting as $\mathbf{w}_{r}$. It is observed that the analytical and simulation results match closely for different BER conditions. The model errors also show a trend of boundness, which demonstrates the rationality of Assumption 6.
Note that model distortion caused by bit errors $\mathbb{E}\left[\delta(w_{d})|\mathbf{w}\right]$ does not necessarily equal zero. For example, any bit error of an all-zero sequence always leads to positive model distortion. Such non-zero bias, which was overlooked in a related analysis [34], is carefully tackled in the convergence analysis in this paper.
After understanding the effect of BER, we then analyze the convergence property of wireless FL with communication error, for which, the following boundness assumption on the global model parameters [14] is needed. The rationality of this assumption can be evident in Fig. 4. From this figure, it can be observed that the model error is bounded under various BER conditions, and thus $M_{\mathbf{w}}$ is also bounded according to (7).
Assumption 6.
(Boundness of Model Parameters [14]) For every communication round $r=0,1,\cdots,R-1$, $M_{r}\triangleq\max(\mathbf{w}_{r})-\min(\mathbf{w}_{r})\leq M$.
@@
With Assumptions 1-6 and if the learning rate $\eta$ satisfies
$1-\frac{K+L}{K}L\eta-2L^{2}\eta^{2}E(E-1)\geq 0,$ (8)
@@
$\begin{split}&\frac{1}{RE}\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\leq\frac{2(\mathbb{E}f(\mathbf{w}_{0})-f^{*})}{\eta RE}\\ &+\frac{(L+1)^{2}DM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}\eta RE}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}\\ &+\frac{L(L+1)\eta(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}+\frac{L^{2}\eta^{2}\sigma_{L}^{2}(K+1)(E-1)}{K},\end{split}$ (9)
where $f^{*}$ is the minimum global training loss, $M$ is defined in Assumption 6, $b_{r}$ is the BER in the $r$-th communication round, and $\mathbf{w}_{0}$ is the initial global model.
@@
Please refer to Appendix B. ∎
@@
It is worthy noting that the first term on the RHS of (9) captures the distance between the initial and minimum global training loss. The second term indicates how bit errors in each communication round affect convergence. The last two terms stem from non-IID data among clients and the random mini-batch sampling.
We next simplify (9) by applying $\eta=\frac{1}{L\sqrt{RE}}=\frac{1}{L\sqrt{T}}$, which can satisfy (8) with a sufficiently large value of $R$.
@@
With Assumptions 1-6 and a sufficiently large value of $R$ such that $\eta=\frac{1}{L\sqrt{RE}}$ satisfies (8), (9) can be simplified as follows:
$\begin{split}\frac{1}{T}&\sum_{r=0}^{R-1}\sum_{e=0}^{E-1}\mathbb{E}\|\nabla f(\mathbf{\bar{w}}_{r,e})\|^{2}\\ \leq&\frac{C_{0}}{\sqrt{T}}+\frac{C_{1}\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}+\frac{C_{2}}{\sqrt{T}}+\frac{C_{3}}{T},\end{split}$ (10)
where $C_{0}\triangleq 2L(\mathbb{E}f(\mathbf{w}_{0})-f^{*})$, $C_{1}\triangleq\frac{L(L+1)^{2}DM^{2}(4^{N}-1)}{3(2^{N}-1)^{2}}$, $C_{2}\triangleq\frac{(L+1)(\sigma_{L}^{2}+\sigma_{G}^{2})}{K}$, $C_{3}\triangleq\frac{L^{2}\sigma_{L}^{2}(K+1)(E-1)}{K}$.
According to Theorem 1, in order to guarantee the convergence rate of $\mathcal{O}(1/\sqrt{T})$, it is crucial to control the BER in each communication round such that the RHS of (10) reduces to zero at a rate of $\mathcal{O}(1/\sqrt{T})$. Since the explicit relationship between $b_{r}$ and $r$ is difficult to obtain, we propose a simple yet practical BER regulation method in Corollary 1. As such, wireless FL systems can achieve the same convergence rate as those with error-free communication.
@@
the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, i.e., wireless FL converges to a local optimal solution of the global training loss function at a rate of $\mathcal{O}(1/\sqrt{T})$.
@@
Denote $b_{r}\propto\frac{1}{(r+1)^{2}}\triangleq\frac{\vartheta}{(r+1)^{2}}$. To prove the RHS of (10) diminishes to zero at a rate of $\mathcal{O}(1/\sqrt{T})$, it is sufficient to verify
$\lim_{T\rightarrow\infty}\frac{\sum_{r=0}^{R-1}b_{r}(1-b_{r})^{N-1}}{\sqrt{T}}=0.$ (12)
Specifically, since
@@
Corollary 1 indicates that gradually decreasing the BER at a rate of $\mathcal{O}(1/(r+1)^{2})$ can guarantee wireless FL converges at a rate of $\mathcal{O}(1/\sqrt{T})$. In other words, the proposed BER regulation rule achieves an improving communication quality over communication rounds. This is consistent with the solution developed in [29] for wireless FL on strongly convex loss functions, which improves the communication quality by increasing the signal-to-noise ratio (SNR) with communication rounds.
III-B BER vs. Number of LDPC Decoding Iterations
The analysis in Section III-A indicates that the convergence rate of conventional FL with error-free communication can still be achieved in wireless FL if the BER is properly controlled. In this subsection, we analyze the relationship between BER and the maximum number of LDPC decoding iterations, based on which, we develop an adaptive decoding scheme to maintain the desired convergence rate.
1 Input: The initial global model $\mathbf{w}_{0}$, $R>1$, $b_{0}$, and $b_{R-1}$.
@@
3       The server broadcasts the $N$-bit representation of the global model $\mathbf{w}_{r}$ and scalars $\min(\mathbf{w}_{r}),\max(\mathbf{w}_{r})$.
4       for Client $k=1,\cdots,K$ do
5             Compute $b_{r}=\frac{(b_{0}-b_{R-1})R^{2}}{(R^{2}-1)(r+1)^{2}}+\frac{b_{R-1}R^{2}-b_{0}}{R^{2}-1}$.
6             Find $Q_{r}$ by its mapping with $b_{r}$.
7             Initialize $decoding.iter=0$.
8             for $decoding.iter<Q_{r}$ do
9                  Execute one iteration of the LDPC decoding algorithm for each frame that contains the global model parameters.
10                   if Parity check passes then
11                        Early termination and break.
@@
15            Reconstruct $\mathbf{\tilde{w}}_{r}^{k}$ according to the LDPC decoding results and $\max(\mathbf{w}_{r}),\min(\mathbf{w}_{r})$.
16             Set $\mathbf{w}^{k}_{r,0}=\mathbf{\tilde{w}}_{r}^{k}$.
@@ 18                   $\mathbf{w}^{k}_{r,e+1}=\mathbf{w}^{k}_{r,e}-\eta\nabla f_{k}(\mathbf{w}^{k}_{r,e};\xi^{k}_{r,e})$
20            Upload $\Delta\mathbf{w}_{r}^{k}=\mathbf{w}^{k}_{r,E}-\mathbf{w}_{r,0}^{k}$ to the server.
@@
22      Perform model aggregation at the server as $\mathbf{w}_{r+1}=\mathbf{w}_{r}+\frac{1}{K}\sum^{K}_{k=1}\Delta\mathbf{w}_{r}^{k}$.
@@
Output: $\mathbf{\bar{w}}_{R,E}=\frac{1}{K}\sum_{k=1}^{K}\mathbf{w}_{R,E}^{k}$ as the learned model.
Algorithm 1 Wireless FL with Adaptive LDPC Decoding
It has been shown in [35] that for message-passing-based channel decoders, to achieve a fraction of 1-$\epsilon$ of the channel capacity over a binary erasure channel, the relationship between the decoding complexity (denoted as $\chi_{D}$) and BER (denoted as $b$) can be modeled as follows:
@@
In the next subsection, we develop an adaptive scheme to control the maximum number of LDPC decoding iterations for clients in a wireless FL system based on Corollary 1 and (16).
III-C LDPC Decoding Iteration Control